OpenStudy (anonymous):

Select the expression equal to ³√250. 2³√125 125³√2 2³√5 5³√2

OpenStudy (raffle_snaffle):

What do you thinK?

OpenStudy (cwrw238):

hint: find a perfect cube which is a factor of 250

OpenStudy (anonymous):

@raffle_snaffle @cwrw238 Is it #1?

OpenStudy (cwrw238):

³√250. = ³√( 2 * 125)

OpenStudy (anonymous):

Is it B? I don't completely understand.

OpenStudy (cwrw238):

³√( 2 * 125) = ³√ 2 * ³√ 125

OpenStudy (cwrw238):

whats the cube root of 125?

OpenStudy (anonymous):

@cwrw238 5

OpenStudy (kc_kennylau):

Factors of 250: 1,2,5,10,25,50,125,250 Perfect cubes: 1,8,27,64,125,216,343 Factors of 250 that is also perfect cubes: 1,125

OpenStudy (kc_kennylau):

Therefore we need to express 250 as 2x125

OpenStudy (kc_kennylau):

\[\Large\sqrt[3]{250}=\sqrt[3]{2\times125}=\sqrt[3]2\times\sqrt[3]{125}=5\sqrt[3]2\]

OpenStudy (cwrw238):

thats it

OpenStudy (anonymous):

Oh okay, I get it now. Thanks @kc_kennylau and @cwrw238 :)

OpenStudy (kc_kennylau):

no problem at all :)